# Rethinking the handling of method failure in comparison studies

### v2

@@ 3.1.3 Check if data matches scope of study
If, given the scope of the study, certain data set characteristics are deemed unrealistic for the targeted readers, authors may consider making sensible parameter changes in the DGM to generate data that is more realistic, in line with White et al., (2024). Similarly, “too extreme” data may also affect real data-based studies. If one of the (real) benchmarking data sets is considered to be irrelevant to the scope of the study after careful examination, authors may consider removing the data set from the study completely. See Section 4.2 for an illustration. However, if the data sets causing failure are deemed relevant to the (simulation or benchmark) study’s scope, the researchers conducting the comparison study should—more appropriate than discarding or imputing in most cases—proceed by considering how real-world users are likely to behave in this situation. In particular, they may modify certain parameters of the concerned method(s) (see Section 3.1.4) or resort to a fallback (see Section 3.1.5). To be able to assess whether data structures associated with failure are too extreme, certain considerations should already be made in the planning stage. They should be aware of the scope and goal of their study: is it the goal to examine the general behavior of the method(s) (i.e., earlier stages of research) or to provide concrete recommendations for real-world practitioners (later stages of research)? Is the focus on testing the methods under regular or more extreme circumstances? If aiming to provide recommendations to real-world users, what degree of extremity in the data remains realistic for them?
@@ 3.1.4 Be realistic when modifying method implementations
In earlier stages of research, it may make sense to modify a method’s implementation entirely (i.e., for all simulation repetitions/benchmark data sets) when encountering method failure in the execution stage. For instance, authors may conclude after execution that the default number of maximum likelihood iterations is generally too low, or that the given optimization algorithm may fail frequently, and that an alternative optimizer is more reliable. This way, authors may modify a greater variety of parameters to see when the method is successful and when it fails. Although this early-stage research is not yet directly aimed at real-world users, authors should remember that it will be used by them later. Therefore, certain practical limitations on parameter modifications should apply to ensure that the method remains feasible for future users.
@@ 3.1.6 Special case: Runtime issues
In contrast to non-convergence and memory issues, method failure is actively declared by the authors when runtime issues occur in the execution stage. That is, running the methods under investigation on hundreds or thousands of data sets requires time restrictions, which are set by the authors arbitrarily, and without which the method would likely produce a valid output. On the other hand, it can sometimes be safely said that real (targeted) users, running the method on a single data set only, are likely to obtain an output. In these cases, when method failure leads to missing rather than undefined performance values, it may make sense to run the experiment on a randomly selected subset of all simulation repetitions/benchmarking data sets for the method affected by runtime issues (“failure by design”). To ensure that this “failure” is independent of data characteristics, it is important to select the subset of data independently of the data sets for which the method has particularly long runtimes. Researchers should therefore think about possible runtime problems and carry out the corresponding random selection of data sets in already the planning stage.
@@ 3.1.7 Report failure proportions, underlying interplays, applied approaches, and share code
Authors of comparison studies should report any occurrences and proportions of method failure. This can serve as an important criterion for real-world users selecting a suitable method. Even the “best performing” method is of limited use if the amount of settings in which it can be applied without frequent failures is limited. Ideally, the proportion of method failure should be included as an additional performance measure for all considered methods (see Morris et al., (2019)); see Zapf et al., (2024) for an example.
Additionally, the interplay of method and data causing failure should be reported. Adding to the previous paragraph, a corresponding reporting for each method can look like “method A, with a failure proportion of $X\%$, fails when a predictor variable is constant across all observations and when a binary predictor has highly imbalanced classes.” In addition to reporting the corresponding failure proportions, readers can learn in which settings a method is applicable, and no less importantly, in which it is not. Additionally, these insights serve as justification of the chosen handlings to method failure. An example of a corresponding reporting can be found in Hornung et al., (2024), who provide detailed descriptions in the supplement of their paper. Authors of comparison studies should additionally report their applied handlings. Any actions resulting from the above-described in-depth investigations and recommendations regarding method failure should be reported with sufficient detail. This includes removing any misused methods, removing data (or modification of DGMs) that do not match the scope of the study, any changes in implementations, and incorporating fallback strategies. Additionally, authors should justify their choices given the scope of the study (i.e., which users it is aimed at), and the research stage in which it takes place.
The complexity of the occurrence and handling of method failure, which becomes clear from the previous sections, emphasizes all the more the value of sharing code, data, and intermediate results. Allowing authors of other, “new” comparison studies in-depth investigations to the specific instances of method failure in the given study may help them to a greater understanding of this topic–and enhance fruitful discussions on suitable handlings.
@@ 4 Illustrations
Note that our focus is exclusively on the handling of undefined performance values and its consequences on the method comparison rather than the method comparison itself. This means that both illustrations represent only a small excerpt (e.g., a few or a single scenario) of what would be a typical publishable study, exclusively focusing on those aspects of the studies that are relevant to the issue of method failure. Importantly, our goal is not to provide insights into the actual performance of the compared methods.
@@ 4.1.1 General
This fictive statistical simulation study deals with the comparison of odds ratio estimation methods. Outcome and exposure are binary ($X,Y\in\left\{0,1\right\}$) and the observed number of subjects with $X=i$ and $Y=j$ is denoted as $n_{ij}$. The scenarios are defined based on different sample sizes $n.obs$, underlying true OR values, and probabilities of exposure $p_{x}$. Especially when the sample size is low, the underlying true OR is far from $1$, or $p_{x}$ is close to 00 or $1$, some simulated data sets may not contain any subjects for one or more of the four outcome-exposure combinations ($n_{ij}=0$). These “sampling zeros” cause calculation issues (e.g., division by 00) for many OR estimation methods, leading to either returning an error or, depending on the position of the sampling zero in the contingency table, non-meaningful OR estimations of 00 or $\infty$” for the corresponding data sets. In practice, users often apply the “Haldane-Anscombe correction” (Haldane, , 1956; Anscombe, , 1956) to handle sampling zeros in the OR estimation. This correction consists of adding an offset of $+0.5$ to all four cells in the contingency table to avoid issues in computation.
@@ 4.1.2 Design of the fictive study
We vary the underlying true OR (OR $\in\{2,3,4,5\}$) and exposure probability $p_{x}$ ($p_{x}\in\{0.25,0.5\})$ while keeping the sample size fixed at $n.obs=50$. This results in eight distinct simulation scenarios (see Table 4 for an overview). For each scenario, we simulate 100 000100000100\,000100 000 outcome-exposure data sets. Five OR estimation methods are applied to the simulated data to obtain a point estimate for the true OR. “Manual” corresponds to the manual and straightforward computation $\widehat{OR}=\frac{n_{11}\cdot n_{00}}{n_{01}\cdot n_{10}}$. The remaining four methods, namely “Fisher,” “Midp,” “Small” (all three implemented in R package epitools (Aragon et al., , 2020)), and “Woolf” (Woolf et al., , 1955) (implemented in R package pairwiseCI (Schaarschmidt and Gerhard, , 2019)) employ more complex modeling strategies. From the five estimators under investigation, only Small and Woolf can handle sampling zeros internally. For each simulation scenario, the estimators are compared regarding bias on the $\log$-scale ($\log$-transformation ensuring symmetry around the value of “no association”).
Our analysis consists of two parts. First, we compare the performance of the OR estimation methods when discarding data sets with sampling zeros for all methods versus only for the failing methods (“ad hoc approaches”; Section 4.1.3). Though users often apply the Haldane-Anscombe correction when encountering sampling zeros, this first part illustrates how the preference of some methods may mainly be driven by the way data sets with sampling zeros are discarded.
Second, we repeat the analysis implementing the principle of fallback strategies suggested in Section 3.1.5. This includes important considerations that must be made when choosing suitable fallbacks. For instance, for the manual estimation of the OR (method “Manual”), only the Haldane-Anscombe correction is applied as a fallback, assuming that a user applying the manual computation in the first place, which is very easy to implement, is unlikely to resort to a method that requires notably more effort. On the other hand, for methods Midp and Fisher for data sets with sampling zeros, we (i) either apply the Haldane-Anscombe correction, (ii) use method Small, or (iii) use method Woolf, assuming that corresponding users are more likely to employ fallbacks that are slightly more difficult to apply. Together with Woolf and Small, this leads to nine “methods” (“pure” methods or pipelines) being compared in total. The results are presented in Section 4.1.4.
@@ 4.1.3 Occurrence of method failure and ad hoc handlings
Table 4 displays the proportion of the 100 000100000100\,000100 000 simulated data sets containing at least one sampling zero across the eight simulation scenarios. As expected, this proportion increases with an increasing true OR$>1$, while decreasing when exposure $X$ is more balanced.
Table 4: Overview of the proportion of the 100 000100000100\,000100 000 simulated data sets with sampling zeros across the simulation scenarios. $n.obs$: number of observations, $p_{x}$: exposure probability.
Simulation scenario n.obs true OR $p_{x}$ Sampling zero proportion
1 50 2 $0.25$ 1.28%
2 50 3 $0.25$ 3.96%
3 50 4 $0.25$ 7.8%
4 50 5 $0.25$ 12.0%
5 50 2 $0.5$ 0.01%
6 50 3 $0.5$ 0.11%
7 50 4 $0.5$ 0.50%
8 50 5 $0.5$ 1.27%
Figure 1: Performance ranks of the OR estimation methods across the eight simulation scenarios based on empirical bias on the $\log$-scale, where a rank of 1 means that the method has the lowest empirical bias in the given scenario. (A) Ranks are compared between discarding data sets for the individual failing methods only (“Single” on the $x$-axis) versus for all methods (“All” on the $x$-axis). (B) The ranks are compared after applying fallback strategies to methods Midp, Fisher, and Manual. For Midp and Small, the fallbacks Small, Woolf, and the Haldane-Anscombe correction (“$+0.5$”) are used. For Manual, only the Haldane-Anscombe correction is used as fallback. Small and Woolf can deal with sampling zeros internally.
Abbreviation of methods: “Fis”: Fisher; “Mid”: Midp“; “Sma”: Small; “Woo”: Woolf; “Man”: Manual (i.e., manual estimation of the OR). (B): The combination of original method $m_{1}$ and fallback $m_{2}$ is marked by “$m_{1}/m_{2}$.”
Ad hoc approaches  For the first analysis, Figure 1(A) shows that the way of discarding data sets with sampling zeros can substantially influence the results of the method comparison. This can lead to changes in the performance rank by up to three positions in some scenarios (see method Woolf) and particularly affect which method is chosen as the best. Note that a more detailed investigation of Woolf’s performance shows that for the data sets without sampling zeros, Woolf slightly underestimates the true OR, while for the few data sets with sampling zeros, Woolf notably overestimates it. This leads to top performance when all data sets are included in its performance assessment (while discarding data sets with sampling zeros for the failing methods).
Generally, the manner of discarding data sets has a larger impact on the performance comparison for scenarios with a higher proportion of sampling zeros. However, differences in rankings within a given scenario are already observable with a proportion of $0.5\%$ of data sets with sampling zeros.
@@ 4.1.4 Recommended handling: using fallbacks
Woolf’s overall good performance across all scenarios is underpinned when using fallbacks for all methods affected by method failure. Furthermore, the comparison between Woolf and Midp, another method performing well in most scenarios, becomes clearer through the use of fallbacks. While they often alternate in rank (and especially between rank 1 and 2) when removing data sets (Figure 1 (A)), their performance comparison is constant across all scenarios when a fallback is applied to Midp (Figure 1 (B)). Since this applies to both fallbacks, the Haldane-Anscombe correction and Small, it is debatable which fallback is preferable. Both fallbacks for Midp frequently switch ranking positions across the performance measures and simulation scenarios, however, fallback Small might be even easier to implement than the Haldane-Anscombe correction, requiring only a single parameter modification from Midp.
@@ 4.2.1 General
The fictive study aims to compare the performance of two methods for constructing confidence intervals (CIs) for the generalization performance of a prediction model based on resampling-based estimates, such as cross-validation. The “naive” method “$N$” for constructing the CIs ignores the dependence between the generalization performances in the resampling repetitions, while method “$C$” corrects for it. The performance of both methods is assessed in terms of coverage. Note that a comprehensive benchmark study of methods for constructing CIs for the generalization error is performed by Schulz-Kümpel et al., (2024).
@@ 4.2.4 Recommended handling
This illustrates well that the further the investigations are taken, the apparent adequate handling of method failure can change notably. In this case, based on the information gained, the adequate handling depends on the aim of the study and requires further considerations such as the following. A prediction problem where splitting often does not occur can be viewed as an edge case. If the study’s scope includes such edge cases, one could attempt to suitably modify the parameters of the classification tree to favor splitting. If, however, the study aims to compare both methods in regular scenarios, one might consider adding additional predictors to favor splitting more strongly, resort to an alternative prediction model (e.g., logistic regression), an alternative measure of generalization performance (e.g., accuracy), or use a different benchmarking data set.
@@ 5 Discussion
Even if we argue for decisions on a case-by-case basis after careful inspection of the failures, we show and illustrate that some general principles hold in most situations. In particular, method failure usually leads to non-existing values instead of just “missing,” i.e., existent but unobserved values. This means imputation, a common but missing-data-inspired approach to handling method failure, is usually inadequate. The same applies to the popular and seemingly logical approach of discarding data associated with failure (either for all or failing methods), which almost never enables the comparison of unconditional performance. In this work, we provide alternative approaches that might be more appropriate to most occurrences of method failure.
An approach that in our view deserves more attention than it receives currently in the literature consists of evaluating whole “pipelines,” including one or several fallback strategies in case of method failure. This approach reflects the typical behavior of most users of methods in practice while enabling the comparison of unconditional performance if constructed carefully: if one method does not work, they resort to an alternative one. Future work is required to collect experiences and define standards regarding such pipelines and their evaluation. Above all, transparent reporting of the frequency of failure and of the corresponding handling is always recommended.
A limitation of our recommendations is that, beyond obvious cases, they typically require both high time resources and thorough expertise in the underlying modelling processes and methods. Researchers conducting comparison studies may not always have these resources to the full extent. This particularly applies to a nuanced examination of the interactions between methods and data sets leading to method failure in different instances. We understand that this bottleneck may impair the full implementation of our recommendations in practice, hence our attempt to provide a slimmed-down version of the recommendations. When time resources and expertise are restricted, transparent reporting is all the more important. Furthermore, this underlines the necessity for openly sharing code, data, and ideally intermediate results, as this allows other researchers to explore the observed instances of method failure in greater detail.
Also, our viewpoint on method failure as leading to “undefined” performance values does not always hold unrestrictedly, which can complicate the implementation of our recommendations. However, with method failure being a complex matter in and of itself, this underscores the importance of a case-by-case evaluation, as described above.
@@
To sum up, we believe there is not “one” right way to handle method failure in all situations. Instead, it is important to carefully choose an approach depending on the context, and then report, explain (and ideally discuss) the decisions made. With our work, we hope to contribute meaningful ideas on how to think about method failure. They lead to helpful strategies for dealing with it, as outlined as part of our recommendations. However, we also believe that method failure is only one of many issues related to comparison studies that deserve much more time, efforts, and methodological strength towards more reliable evidence on the behavior of methods in the context of a “replication crisis in methodological research” (Boulesteix et al., , 2020). We need more well-designed neutral comparison studies (Boulesteix et al., , 2017)—late phase studies according to the terminology of Heinze et al., (2024). In such studies, no time is spent on the development of new methods. This leaves time for implementing our recommendations and diving into the technical details of the methods under investigation.

### v3

@@ 3.1.3 Check if data matches scope of study
If, given the scope of the study, certain data set characteristics are deemed unrealistic for the targeted readers, authors may consider making sensible parameter changes in the DGM to generate data that is more realistic, in line with White et al., (2024). Similarly, “too extreme” data may also affect real data-based studies. If one of the (real) benchmark data sets is considered to be irrelevant to the scope of the study after careful examination, authors may consider removing the data set from the study completely. See Section 4.2 for an illustration. However, if the data sets causing failure are deemed relevant to the (simulation or benchmark) study’s scope, the researchers conducting the comparison study should—more appropriate than discarding or imputing in most cases—proceed by considering how real-world users are likely to behave in this situation. In particular, they may modify certain parameters of the concerned method(s) (see Section 3.1.4) or resort to a fallback (see Section 3.1.5). To be able to assess whether data structures associated with failure are too extreme, certain considerations should already be made in the planning stage. Researchers should be aware of the scope and goal of their study: is it the goal to examine the general behavior of the method(s) (i.e., earlier stages of research) or to provide concrete recommendations for real-world practitioners (later stages of research)? Is the focus on testing the methods under regular or more extreme circumstances? If aiming to provide recommendations to real-world users, what degree of extremity in the data remains realistic for them?
@@ 3.1.4 Be realistic when modifying method implementations
In earlier stages of research, it may make sense to modify a method’s implementation entirely (i.e., for all simulation repetitions/benchmark data sets) when encountering method failure during execution. For instance, authors may conclude after execution that the default number of maximum likelihood iterations is generally too low, or that the given optimization algorithm may fail frequently, and that an alternative optimizer is more reliable. This way, authors may modify a greater variety of parameters to see when the method is successful and when it fails. Although this early-stage research is not yet directly aimed at real-world users, authors should remember that it will be used by them later. Therefore, certain practical limitations on parameter modifications should apply to ensure that the method remains feasible for future users.
@@ 3.1.6 Special case: Runtime issues
In contrast to non-convergence and memory issues, method failure is actively declared by the authors when runtime issues occur in the execution stage. That is, running the methods under investigation on hundreds or thousands of data sets requires time restrictions, which are set by the authors arbitrarily, and without which the method would likely produce a valid output. On the other hand, it can sometimes be safely said that real (targeted) users, running the method on a single data set only, are likely to obtain an output. In these cases, when method failure leads to missing rather than undefined performance values, it may make sense to run the experiment on a randomly selected subset of all simulation repetitions/benchmark data sets for the method affected by runtime issues (“failure by design”). To ensure that this “failure” is independent of data characteristics, it is important to select the subset of data independently of the data sets for which the method has particularly long runtimes. Researchers should therefore think about possible runtime problems and carry out the corresponding random selection of data sets already in the planning stage.
@@ 3.1.7 Report failure proportions, underlying interplays, applied approaches, and share code
Authors of comparison studies should report any occurrences and proportions of method failure. This can serve as an important criterion for real-world users selecting a suitable method. Even the “best performing” method is of limited use if the number of settings in which it can be applied without frequent failures is limited. Ideally, the proportion of method failure should be included as an additional performance measure for all considered methods (see Morris et al., (2019)); see Zapf et al., (2024) for an example.
Additionally, the interplay of method and data causing failure should be reported. Adding to the previous paragraph, a corresponding reporting for each method can look like “method A, with a failure proportion of $X\%$, fails when a predictor variable is constant across all observations and when a binary predictor has highly imbalanced classes.” In addition to reporting the corresponding failure proportions, readers can learn in which settings a method is applicable, and no less importantly, in which it is not. Additionally, these insights serve as justification for the chosen handlings of method failure. An example of a corresponding reporting can be found in Hornung et al., (2024), who provide detailed descriptions in the supplement of their paper.
Authors of comparison studies should additionally report their applied handlings with sufficient detail, that is, any actions resulting from the above-described in-depth investigations and recommendations regarding method failure. This includes removing any misused methods, removing data (or modifying DGMs) that do not match the scope of the study, any changes in implementations, and incorporating fallback strategies. Additionally, authors should justify their choices given the scope of the study (i.e., which users it is aimed at) and the research stage in which it takes place.
The complexity of the occurrence and handling of method failure, which becomes clear from the previous sections, emphasizes all the more the value of sharing code, data, and intermediate results. Allowing authors of other, “new” comparison studies in-depth investigations of the specific instances of method failure in the given study may help them gain a greater understanding of this topic–and enhance fruitful discussions on suitable handlings.
@@ 4 Illustrations
Note that our focus is exclusively on the handling of method failure and its consequences on the method comparison rather than the method comparison itself. This means that both illustrations represent only a small excerpt (e.g., a few or a single scenario) of what would be a typical publishable study, exclusively focusing on those aspects of the studies that are relevant to the issue of method failure. Importantly, our goal is not to provide insights into the actual performance of the compared methods.
@@ 4.1.1 General
This fictive statistical simulation study deals with the comparison of odds ratio (OR) estimation methods. Outcome and exposure are binary ($X,Y\in\left\{0,1\right\}$) and the observed number of subjects with $X=i$ and $Y=j$ is denoted as $n_{ij}$. The simulation scenarios are defined by the sample size $n.obs$, the underlying true OR, and the probability of exposure $p_{x}$. Especially when the sample size is low, the underlying true OR is far from $1$, or $p_{x}$ is close to 0 or $1$, some simulated data sets may not contain any subjects for one or more of the four outcome-exposure combinations ($n_{ij}=0$). These “sampling zeros” cause calculation issues (e.g., division by 0) for many OR estimation methods, leading to either returning an error or, depending on the position of the sampling zero in the contingency table, non-meaningful OR estimations of 0 or $\infty$” for the corresponding data sets. In practice, users often apply the “Haldane-Anscombe correction” (Haldane,, 1956; Anscombe,, 1956) to handle sampling zeros in the OR estimation. This correction consists of adding an offset of +0.5+0.5 to all four cells in the contingency table to avoid issues in computation.
@@ 4.1.2 Design of the fictive study
We vary the underlying true OR (OR{2,5}OR\ $\in\{$2,5\}) and exposure probability $p_{x}$ ($p_{x}\in\{0.25,0.5\})$ while keeping the sample size fixed at the relatively low value of $n.obs=50$. This results in four distinct simulation scenarios (see Table 5 for an overview). For each scenario, we simulate 100 000100\,000 outcome-exposure data sets. Five OR estimation methods are applied to the simulated data to obtain a point estimate for the true OR. “Manual” corresponds to the manual and straightforward computation $\widehat{OR}=\frac{n_{11}\cdot n_{00}}{n_{01}\cdot n_{10}}$. The remaining four methods, namely “Fisher,” “Midp,” “Small” (all three implemented in R package epitools (Aragon et al.,, 2020)), and “Woolf” (Woolf et al.,, 1955) (implemented in R package pairwiseCI (Schaarschmidt and Gerhard,, 2019)) employ more complex modeling strategies. From the five estimators under investigation, only Small and Woolf can handle sampling zeros internally. For each simulation scenario, the estimators are compared regarding bias on the $\log$-scale ($\log$-transformation ensuring symmetry around the value of “no association”).
Our analysis consists of two parts. First, we compare the performance of the OR estimation methods when discarding data sets with sampling zeros for all methods versus only for the failing methods (“ad hoc approaches”; Section 4.1.3). Though real users often apply the Haldane-Anscombe correction when encountering sampling zeros, this first part illustrates how the preference of some methods may mainly be driven by the way data sets with sampling zeros are discarded.
Second, we repeat the analysis implementing the principle of fallback strategies suggested in Section 3.1.5. This includes important considerations that must be made when choosing suitable fallbacks. For instance, for the manual estimation of the OR (method “Manual”), only the Haldane-Anscombe correction is applied as a fallback, assuming that a user applying the manual computation in the first place, which is very easy to implement, is unlikely to resort to a method that requires notably more effort. On the other hand, for methods Midp and Fisher for data sets with sampling zeros, we either (i) apply the Haldane-Anscombe correction, (ii) use method Small, or (iii) use method Woolf, assuming that corresponding users may also employ fallbacks that are slightly more difficult to apply. Together with Woolf and Small, this leads to nine “methods” (“pure” methods or pipelines) being compared in total. The results are presented in Section 4.1.4.
@@ 4.1.3 Occurrence of method failure and ad hoc handlings
Table 5 displays the proportion of the 100 000100\,000 simulated data sets containing at least one sampling zero across the four simulation scenarios. As expected, this proportion increases with an increasing true OR>1>1, while decreasing when exposure $X$ is more balanced.
Table 5: Overview of the proportion of the 100 000100\,000 simulated data sets (fixed sample size $n.obs=50$) with sampling zeros across the simulation scenarios. $p_{x}$: exposure probability.
Simulation scenario true OR $p_{x}$ Sampling zero proportion
1 2 $0.25$ 1.28%
2 5 $0.25$ 12.0%
3 2 $0.5$ 0.01%
4 5 $0.5$ 1.27%
Figure 1: Performance ranks of the OR estimation methods across the eight simulation scenarios based on empirical bias on the $\log$-scale, where a rank of 1 means that the method has the lowest empirical bias in the given scenario. (A) Ranks are compared between discarding data sets for the individual failing methods only (“Single” on the $x$-axis) versus for all methods (“All” on the $x$-axis). (B) The ranks are compared after applying fallback strategies to methods Midp, Fisher, and Manual. For Midp and Small, the fallbacks Small, Woolf, and the Haldane-Anscombe correction (“+0.5+0.5”) are used. For Manual, only the Haldane-Anscombe correction is used as fallback. Small and Woolf can deal with sampling zeros internally.
Abbreviation of methods: “Fis”: Fisher; “Mid”: Midp“; “Sma”: Small; “Woo”: Woolf; “Man”: Manual (i.e., manual estimation of the OR). (B): The combination of original method $m_{1}$ and fallback $m_{2}$ is marked by $m_{1}/\penalty 50m_{2}$.”
Ad hoc approaches In the first analysis, Figure 1(A) demonstrates that the way of discarding data sets with sampling zeros can substantially affect the results of the method comparison, with discrepancies regarding performance ranks already observable at a sampling zero proportion of just over $1\%$. This impact becomes more pronounced as the proportion of sampling zeros increases, making a difference in the observed performance by up to three ranks. In particular, it can affect which method is perceived as best-performing. Note that a more detailed investigation of Woolf’s performance shows that for the data sets without sampling zeros, Woolf slightly underestimates the true OR, while for the few data sets with sampling zeros, Woolf notably overestimates it. This leads to top performance when all data sets are included in its performance assessment (while discarding data sets with sampling zeros for the failing methods).
@@ 4.1.4 Recommended handling: using fallbacks
Woolf’s overall good performance across all scenarios is underpinned when using fallbacks for all methods affected by method failure. Furthermore, the comparison between Woolf and Midp, another method performing well in most scenarios, becomes clearer through the use of fallbacks. While they often alternate in rank (and especially between rank 1 and 2) when removing data sets (Figure 1 (A)), their performance comparison is constant across all scenarios when a fallback is applied to Midp (Figure 1 (B)). Since this applies to both fallbacks, the Haldane-Anscombe correction and Small, it is debatable which fallback is preferable. Both fallbacks for Midp frequently switch ranking positions across the performance measures and simulation scenarios, however, fallback Small might be even easier to implement than the Haldane-Anscombe correction, requiring only a simple code modification from Midp.
@@ 4.2.1 General
The fictive study aims to compare the performance of two methods for constructing confidence intervals (CIs) for the generalization performance of a prediction model based on resampling-based estimates, such as cross-validation. The “naive” method “NN” for constructing the CIs ignores the dependence between the generalization performances in the resampling repetitions, while method “CC” corrects for it. The performance of both methods is assessed in terms of coverage. Note that a comprehensive benchmark study of methods for constructing CIs for the generalization error is performed by Schulz-Kümpel et al., (2024).
@@ 4.2.4 Recommended handling
This illustrates well that the further the investigations are taken, the apparent adequate handling of method failure can change notably. In this case, based on the information gained, adequate handling depends on the aim of the study and requires further considerations such as the following. A prediction problem where splitting often does not occur can be viewed as an edge case. If the study’s scope includes such edge cases, one could attempt to suitably modify the parameters of the classification tree to favor splitting. If, however, the study aims to compare both methods in regular scenarios, one might consider adding additional predictors to favor splitting more strongly, resort to an alternative prediction model (e.g., logistic regression), an alternative measure of generalization performance (e.g., accuracy), or use a different benchmark data set.
@@ 5 Discussion
Even if we argue for decisions on a case-by-case basis after careful inspection of the failures, we show and illustrate that some general principles hold in most situations. In particular, method failure usually leads to non-existing values instead of just “missing,” i.e., existent but unobserved values. This means imputation, a common but missing-data-inspired approach to handling method failure, is usually inadequate. The same applies to the popular and seemingly logical approach of discarding data associated with failure (either for all or the failing methods), which almost never enables the comparison of unconditional performance. In this work, we provide alternative approaches that might be more appropriate to most occurrences of method failure.
An approach that in our view deserves more attention than it receives currently in the literature consists of evaluating whole “pipelines” of methods, including one or several fallback strategies in case of method failure. This approach reflects the typical behavior of most users of methods in practice while enabling the comparison of unconditional performance if constructed carefully: if one method does not work, they resort to an alternative one. Future work is required to collect experiences and define standards regarding such pipelines and their evaluation. Above all, transparent reporting of the frequency of failure and of the corresponding handling is always recommended.
A limitation of our recommendations is that, beyond obvious cases, they typically require both high time resources and thorough expertise in the underlying modelling processes and methods. Researchers conducting comparison studies may not always have these resources to the full extent. This particularly applies to a nuanced examination of the interactions between methods and data sets leading to method failure in different instances. We understand that this bottleneck may impair the full implementation of our recommendations in practice, hence our attempt to provide a slimmed-down version of the recommendations. When time resources and expertise are restricted, transparent reporting is all the more important. Furthermore, this underlines the necessity for openly sharing code, data, and ideally intermediate results, as this allows other researchers to explore the observed instances of method failure in greater detail. Also, our viewpoint on method failure as leading to “undefined” performance values does not always hold unrestrictedly, which can complicate the implementation of our recommendations. However, with method failure being a complex matter in and of itself, this underscores the importance of a case-by-case evaluation, as described above.
@@
To sum up, we believe there is not “one” right way to handle method failure in all situations. Instead, it is important to carefully choose an approach depending on the context, and then report, explain (and ideally discuss) the decisions made. With our work, we hope to contribute meaningful ideas on how to think about method failure. They lead to helpful strategies for dealing with it, as outlined as part of our recommendations. However, we also believe that method failure is only one of many issues related to comparison studies that deserve much more time, effort, and methodological strength towards more reliable evidence on the behavior of methods in the context of a “replication crisis in methodological research” (Boulesteix et al.,, 2020). We need more well-designed neutral comparison studies (Boulesteix et al.,, 2017)—late phase studies according to the terminology of Heinze et al., (2024). In such studies, no time is spent on the development of new methods. This leaves time for implementing our recommendations and diving into the technical details of the methods under investigation.